OpenStudy (anonymous):

USe a method of variation of parameters to find the general solution to the systems of differential equations X'= (-2 5) X + ( tan t ) -1 2 1/(cos t)

OpenStudy (anonymous):

x'=x*A+B(t) x'+x*(-A)=B(t) x'*e^(-At) +x*(-A)*e^(-At)=B(t)*e^(-At) (x*e^(-At))'=B(t)*e^(-At) =>x*e^(-At)=***** ( take integral by part)